# ClassDiffusion: More Aligned Personalization Tuning with Explicit Class Guidance

### v1

@@ Abstract
Recent text-to-image customization works have been proven successful in generating images of given concepts by fine-tuning the diffusion models on a few examples. However, these methods tend to overfit the concepts, resulting in failure to create the concept under multiple conditions (e.g., headphone is missing when generating “a <sks> dog wearing a headphone”). Interestingly, we notice that the base model before fine-tuning exhibits the capability to compose the base concept with other elements (e.g., “a dog wearing a headphone”), implying that the compositional ability only disappears after personalization tuning. Inspired by this observation, we present ClassDiffusion, a simple technique that leverages a semantic preservation loss to explicitly regulate the concept space when learning the new concept. Despite its simplicity, this helps avoid semantic drift when fine-tuning on the target concepts. Extensive qualitative and quantitative experiments demonstrate that the use of semantic preservation loss effectively improves the compositional abilities of the fine-tune models. In response to the ineffective evaluation of CLIP-T metrics, we introduce BLIP2-T metric, a more equitable and effective evaluation metric for this particular domain. We also provide in-depth empirical study and theoretical analysis to better understand the role of the proposed loss. Lastly, we also extend our ClassDiffusion to personalized video generation, demonstrating its flexibility. Our project page are https://classdiffusion.github.io/.
@@ 1 Introduction
Thanks to the rapid progress in the diffusion model [31, 70, 63, 46, 53, 57, 61, 65, 91, 66], the field of text-to-image generation has achieved significant progress in recent years. Leading text-to-image models [64, 20, 37, 1, 78, 39, 33] have been successful in generating high-fidelity images that align well with textual inputs. Recently, a significant portion of research  [4, 90, 77, 89, 86, 58, 2, 5, 11, 28, 35, 8] has transitioned their focus from creating high-quality images to enhancing control over the generated images. Among these works, an important and widely explored research domain is subject-driven personalized generation, which aims to generate new images for specific concept given some reference images of that concept.
@@
Existing personalization methods  [64, 20, 37, 1, 78, 39, 33, 9, 76, 49, 83, 79, 21] can generate images that closely resemble the concept by fine-tuning the base text-to-image model in a specific image set. However, this process leads to a weakening in the model’s compositional ability. For example, when generating “a <sks> dog wearing a headphone”, though the given dog is well reconstructed, the headphone is always missing (Fig. 1). A commonly accepted explanation within the community [64, 73, 29, 20] attributes this phenomenon to overfitting given limited number of images. Nonetheless, the fundamental cause of this overfitting remains unexplored. In this work, we aim to investigate the underlying causes behind the overfitting.
@@
Figure 1: The base Stable Diffusion (SD) possesses the capbility to compose the concept of a dog and headphone, generating a dog wearing a headphone. However, we notice that this compositional generation capability is lost during personalization tuning. For example, when using CustomDiffusion (CD) [37], the headphone is missing despite the target corgi is generated successfully. On the other hand, our method can successfully compose the target corgi with headphone.
@@
Upon initial examination, it appears that the model diminishes some of its original capabilities after personalization tuning. Taking Stable Diffusion (SD) [72] as an example, from Fig. 1, we notice that the base SD model actually has acquired the skill to compose the concept of dog and a headphone. Nevertheless, after fine-tuning, it struggles to achieve similar levels of compositional generation (e.g., headphone is missing although the target concept <sks> dog can be generated successfully). We hypothesized that the decline in this compositional ability stems from the semantic drift of the target concept away from its superclass during fine-tuning. To better understand this, we conduct some empirical analysis by visualizing the text space and cross-attention layers in Fig. 22. In addition, we also perform theoretical analysis and find that the root cause lies in the semantic bias that reduces the entropy of the probability of the composed conditions, which significantly increases the difficulty to simultaneously sample the target concept combined with other elements.
@@
Based on our experimental findings and theoretical analysis, we introduce ClassDiffusion to tackle the issue of weakening compositional capacity post-fine-tuning. This method uses semantic preservation loss to explicitly guide the model to restore the semantic imbalance that arises during the fine-tuning stage. In particular, it narrows the gap between the text embeddings of the target concept and its respective superclass in the textual space. Despite its simplicity, the proposed loss can successfully recover the compositional ability as shown in Fig. 1. Extensive experiments have demonstrated the effectiveness of our method to restore the base model’s compositional generation capability. Furthermore, we explore the potential of our approach in personalized video synthesis, showcasing its flexibility. In addition, we found that CLIP-T metric can hardly reflect the actual performance of personalized generation. We therefore introduce BLIP2-T metric, a more equitable and effective evaluation metric for this particular domain. To summarize, the contributions of our work are:
@@
  • We offer a thorough examination to understand why existing tuning-based subject-driven personalization methods suffer from the loss of compositional ability. This is elucidated through both experimental observations and theoretical analysis.
@@
  • We propose ClassDiffusion, a simple technique to recover the compositional capabilities lost during personalization tuning by adding a simple semantic preservation loss.
  • We proposed BLIP2-T Score, a fairer, better-performing evaluation metric for evaluating image-text alignment.
@@ 2 Related Work
Text-to-image generation is designed to generate high-quality, high-fidelity images that are aligned with textual prompts. This field has been under research for an extended period. Recently, the field of Text-to-image generation has made significant progress with extensive research in Generative Adversarial Network (GAN)  [24, 92, 23, 51], Variational Autoencoder (VAE)  [36, 6, 13, 74], and Diffusion models  [31, 70, 63, 46, 53, 57, 61, 65, 91, 66]. The diffusion models achieve a new state-of-the-art (SOTA) in unconditional image generation. Numerous works [12, 44, 53, 61, 65, 60, 59] have been done to make the image generated by diffusion models more aligned with the textual prompts. Among them, Stable Diffusion  [63] is a widely recognized model in the field, utilizes a cross-attention mechanism to integrate textual conditions into the image generation process and employs the Latent Diffusion Model, which maps the image to latent space [63]. Our research is based on the Stable Diffusion framework due to its adaptability and wide use in the community. Furthermore, different methods exist for controlling generative models. The primary categories for controlling generative models typically include Text-guided  [18, 22, 7, 58, 62], Image-guided  [64, 20, 37, 1, 78, 39, 33], Additional Sparse conditions [4, 90, 77, 89, 86, 58, 2, 5, 11, 28, 35, 8, 47], Brain-guided  [10, 72, 55, 45, 3, 19, 52], Sound-Guided [58, 87], and some universe control [47, 41, 88]. Text-guided control utilizes textual descriptions to directly influence the outcome, guiding the model based on specific verbal instructions. Our method focused on the text-guided controllable generative model.
@@
Subject-driven personalized generation is focused on creating images based on reference images. Recent works [64, 20, 85, 37, 1, 78, 39, 33, 9, 76, 49, 83, 79, 21, 56, 71] have explored techniques for producing striking resemblance images in multiple ways. One of the primary ways is to fine-tune the base text-to-image models. Furthermore, there has been a significant effort in research  [14, 37, 29, 69, 78, 82, 32, 38, 34] aimed at integrating various concepts in personalization. While producing striking resemblance images, fine-tuning the base model on a small set of images leads to overfitting, resulting in unexpected issues. One prevalent issue discussed in prior research is the decrease in diversity. Recent studies have proposed various methods to address these issues. For instance, DreamBooth  [64] introduced the Class-Specific Priori Loss, which effectively addresses diversity reduction by recovering the class’s prior knowledge. Another common issue is the inability to generate images under multiple conditions. Some Recent Research  [39, 30, 73, 48, 29] proposed some methods to migrate this appearance. However, the underlying reasons for this phenomenon have not been thoroughly investigated. Our research endeavors to explore these reasons and develop solutions to overcome this challenge.
Figure 2: (a) Each dot represents the CLIP text embedding of a phrase combining an adjective and “dog” (e.g., a cute dog). After fine-tuning, customized concepts (the blue dot represents the concept before fine-tuning, and the red dot represents the one after) become far away from the center of the “dog” distribution in the text feature space. More detailed analysis can be found in Appendix D. (b) Visualization results of cross-attention maps corresponding to the dog token when using the prompt “a photo of a dog swimming in the pool”.
Figure 3: During the personalization tuning process, as the distribution of dogs shrinks, the conditional distribution of dogs and headphones also shrinks. This gradually increases the difficulty to sample in this distribution, leading to the weakening of the compositional generation capability. Our ClassDiffusion mitigates this by incorporating semantic preservation loss (SPL) to minimize the semantic drift of the personalized concept from its superclass.
@@ 3.1 Preliminary
Stable Diffusion is used by almost all of the methods in the area of subject-driven personalization generation, which is developed from the Latent Diffusion Model (LDM)  [63]. For any input image, Stable Diffusion first transforms it into a latent representation $x$ using the encoder $\varepsilon$ of a Variational Autoencoder  [36]. The diffusion process then operates on $x$ by incrementally introducing noise, resulting in a fixed-length Markov Chain represented as $x_{1},x_{2},...,x_{T}$, where $T$ is the chain’s length. Stable Diffusion utilizes a UNet architecture to learn the reverse of this diffusion process, predicting a denoised version of the latent input $x_{t}$ at each timestep $t$ from 1 to $T$. In the context of text-to-image generation, the text prompts’ conditioning information $y$ is encoded into an intermediate representation $\tau_{\theta}(y)=c$, where $\tau_{\theta}$ is a pre-trained CLIP  [59] text encoder. The primary objective in training this text-to-image diffusion model involves optimizing this transformation and prediction process, and it can be expressed as:
@@ 3.2 Experimental Analysis
We begin by conducting some empirical analysis to investigate the reason behind the lost of compositional ability after personalization tuning.
Semantic drift in CLIP text-space. We first examine the CLIP text sample space before and after fine-tuning. We visualize the CLIP text space in Fig. 2 and notice that customized concepts are more distant from the class center. Specifically, among the 71 phrases composing the super-category (“dog” in this case) with different adjectives (e.g., "a cute dog", "a tall dog" etc.), we observe a significant increase in distance of the target concept from its corresponding superclass. This indicates that the semantics of the personalized concepts have changed during fine-tuning, which potentially leads to an inability to access the knowledge associated with the super-categories learned during pre-training. More details can be found in Appendix D,  E.
Reduction of cross-attention strength. We further investigate the model by visualizing the cross-attention layers in Fig. 2. It shows that, while the "dog" token activates the relevant region in the image, its activation level is notably lower than that of the pre-trained model. These findings align with our observations in the CLIP text space and provide support for our hypothesis regarding this discrepancy. Next, we theoretically analyze why the semantic drift of personalized phrases results in the weakening of the compositional ability from the respective of reduction in the entropy of composable conditional probability.
@@ 3.3 Theoretical Analysis
Figure 4: Overview of our proposed ClassDiffusion. Our semantic preservation loss (SPL) is calculated by measuring the cosine distance between text features extracted from the same text transformer (using EOS tokens as text features following CLIP) for phrases with personalized tokens and phrases with only superclasses.
@@
$p(x|c_{1},c_{2},\dots,c_{i})\propto p(c_{class}|x)p(x)\prod_{i\in T}\frac{p(c_{i})p(x|c_{i})}{p(x)}$ (4)
@@
As a result, it is more difficult to sample from our demanded conditional distributions under $c_{\text{class}},c_{1},...,c_{i}$ conditions than before the fine-tuning, leading to the phenomenon that the combining ability is weakened after the fine-tuning. We will discuss the theoretical reasons here in more detail in Appendix C. Fig.  3 illustrates the changes in the distribution during this process that lead to a weakening of the compositional generation capability.
Inspired by the empirical study and theoretical analysis above, in the next section, we present our proposed semantic preservation loss for mitigating the semantic drift occurred during fine-tuning.

### v3

@@ Abstract
Recent text-to-image customization works have proven successful in generating images of given concepts by fine-tuning the diffusion models on a few examples. However, tunning-based methods inherently tend to overfit the concepts, resulting in failure to create the concept under multiple conditions (e.g., headphone is missing when generating “a <sks> dog wearing a headphone”). Interestingly, we notice that the base model before fine-tuning exhibits the capability to compose the base concept with other elements (e.g., “a dog wearing a headphone”), implying that the compositional ability only disappears after personalization tuning. We observe a semantic shift in the customized concept after fine-tuning, indicating that the personalized concept is not aligned with the original concept, and further show through theoretical analyses that this semantic shift leads to increased difficulty in sampling the joint conditional probability distribution, resulting in the loss of the compositiona ability. Inspired by this finding, we present ClassDiffusion, a technique that leverages a semantic preservation loss to explicitly regulate the concept space when learning the new concept. Although simple, this approach effectively prevents semantic drift during the fine-tuning process on the target concepts. Extensive qualitative and quantitative experiments demonstrate that the use of semantic preservation loss effectively improves the compositional abilities of fine-tuning models. Lastly, we also extend our ClassDiffusion to personalized video generation, demonstrating its flexibility.
@@ 1 Introduction
Thanks to the rapid progress in the diffusion model [31, 72, 65, 48, 55, 59, 63, 67, 95, 68, 92], the field of text-to-image generation has achieved significant progress in recent years. The leading text-to-image models [66, 20, 38, 1, 80, 40, 34] have been successful in generating high-fidelity images that align well with textual inputs. Recently, a significant part of the research  [4, 94, 79, 93, 88, 60, 2, 5, 11, 28, 36, 8] has changed their focus from creating high-quality images to improving control over the generated images. Among these works, an important and widely explored research domain is subject-driven personalized generation, which aims to generate new images for a specific concept given some reference images of that concept.
@@
Existing personalization methods  [66, 20, 38, 1, 80, 40, 34, 9, 78, 51, 85, 81, 21] can generate images that closely resemble the concept by fine-tuning the base text-to-image model in a specific image set. However, all tuning-based models will inherently suffer from the over-fitting introduced by this process, which leads to weakening in the compositional ability of the model. For example, when generating “a <sks> dog wearing a headphone”, though the given dog is well reconstructed, the headphone is always missing (Fig. 1). This feature affects the diversity of the generated output in practical use. A commonly accepted explanation within the community [66, 75, 29, 20] attributes this phenomenon to overfitting given a limited number of images. However, the fundamental cause of this overfit remains unexplored. In this work, our aim is to investigate the underlying causes behind the overfitting.
@@
Figure 1: The base Stable Diffusion (SD) possesses the capbility to compose the concept of a dog and headphone, generating a dog wearing a headphone. However, we notice that this compositional generation capability is lost during personalization tuning. For example, when using Custom Diffusion (CD) [38], the headphone is missing despite the target corgi is generated successfully. On the other hand, our method can successfully compose the target corgi with the headphone.
@@
Upon initial examination, it appears that the model diminishes some of its original capabilities after personalization tuning. Taking Stable Diffusion (SD) [74] as an example, from Fig. 1, we observe that the base SD model indeed has the ability to combine the concepts of a dog and a headphone. However, after fine-tuning, the model struggles to achieve compositional generation; for instance, while the target concept <sks> (dog) can be generated successfully, the headphone is missing. We hypothesized that the decline in this compositional ability stems from the semantic drift of the target concept away from its superclass target during fine-tuning. To better understand this, we conduct some empirical analysis by visualizing the CLIP text-space and cross-attention map activation area in Fig. 33. In addition, we also perform theoretical analysis and find that the root cause lies in the semantic bias that reduces the entropy of the probability of the composed conditions, which significantly increases the difficulty to simultaneously sample the target concept combined with other elements.
@@
Based on our experimental findings and theoretical analysis, we introduce ClassDiffusion to address the issue of weakening compositional capacity after fine-tuning. Fig. 2 shows the performance of our method. Our method uses semantic preservation loss to explicitly guide the model to restore the semantic imbalance that arises during the fine-tuning stage. In particular, it narrows the gap between the text embeddings of the target concept and its respective superclass in the textual space. Despite its simplicity, the proposed loss can successfully recover the compositional ability as shown in Fig. 1. Therefore, distinct from prevalent loss design in the community which seek to migrate the overfitting in tuning-based models, our method enhances the model’s capacity of following the text prompt while maintaining the concept of a customized subject. Extensive experiments have demonstrated the effectiveness of our method in restoring the compositional generation capability of the base model. Furthermore, we explore the potential of our approach in personalized video synthesis, showcasing its ability in recovering the semantical space of the generative model. In addition, we found that the CLIP-T metric can hardly reflect the actual performance of personalized generation. Therefore, we introduce the BLIP2-T metric, a more equitable and effective evaluation metric for this particular domain. To summarize, the contributions of our work are:
@@
  • We offer a thorough examination to understand why existing tuning-based subject-driven personalized methods inherently suffer from the loss of compositional ability. This is elucidated through both experimental observations and theoretical analysis.
@@
  • We propose ClassDiffusion, a simple technique to recover the compositional capabilities lost during personalized tuning.
  • Extensive experiments demonstrate that the proposed technique achieves improved personalization ability in image and video generation tasks.
Figure 2: A qualitative result of two small stories produced by our model. The above showcases a bear’s literary journey: from reading a book to ultimately earning a Nobel Literature Prize. The below shows the fate of a sunglasses. Finally, the bear gets the sunglasses. It shows a potential real-world application due to our model’s high performance.
@@ 2 Related Work
Text-to-image generation is designed to generate high-quality, high-fidelity images that are aligned with textual prompts. This field has been under research for an extended period. Recently, the field of Text-to-image generation has made significant progress with extensive research in Generative Adversarial Network (GAN)  [24, 98, 23, 53], Variational Autoencoder (VAE)  [37, 6, 13, 76], and Diffusion models  [31, 72, 65, 48, 55, 59, 63, 67, 95, 68]. The diffusion models achieve a new state-of-the-art (SOTA) in unconditional image generation. Numerous works [12, 46, 55, 63, 67, 62, 61, 96, 91, 43, 33, 99] have been done to make the image generated by diffusion models more aligned with the textual prompts. Among them, Stable Diffusion  [65] is a widely recognized model in the field, utilizes a cross-attention mechanism to integrate textual conditions into the image generation process and employs the Latent Diffusion Model, which maps the image to latent space [65]. Our research is based on the Stable Diffusion framework due to its adaptability and wide use in the community. Furthermore, different methods exist for controlling generative models. The primary categories for controlling generative models typically include Text-guided  [18, 22, 7, 60, 64], Image-guided  [66, 20, 38, 1, 80, 40, 34], Additional Sparse conditions [4, 94, 79, 93, 88, 60, 2, 5, 11, 28, 36, 8, 49], Brain-guided  [10, 74, 57, 47, 3, 19, 54], Sound-Guided [60, 89], and some universe control [49, 42, 90]. Text-guided control utilizes textual descriptions to directly influence the outcome, guiding the model based on specific verbal instructions. Our method focused on the text-guided controllable generative model.
@@
Subject-driven personalized generation is focused on creating images based on reference images. Recent works [66, 20, 87, 38, 1, 80, 40, 34, 9, 78, 51, 85, 81, 21, 58, 73] have explored techniques for producing striking resemblance images in multiple ways. One of the primary ways is to fine-tune the base text-to-image models. Furthermore, there has been a significant effort in research  [14, 38, 29, 71, 80, 84, 32, 39, 35] aimed at integrating various concepts in personalization. While striking resemblance images are produced, fine-tuning the base model on a small set of images leads to overfitting, resulting in unexpected issues. One prevalent issue discussed in prior research is the decrease in diversity. Recent studies have proposed various methods to address these issues. For instance, DreamBooth  [66] introduced the Class-Specific Priori Loss, which effectively addresses diversity reduction by recovering the class’s prior knowledge. However, it does not effectively maintain the ability of the model to follow the text prompt. Another common issue is the inability to generate images under multiple conditions. Some Recent Research  [40, 30, 75, 50, 29] proposed some methods to migrate this appearance. However, the underlying reasons for this phenomenon have not been thoroughly investigated. Our research endeavors to explore these reasons and develop solutions to overcome this challenge.
@@ 3.1 Preliminary
Stable Diffusion [65] is widely used in image generation task. For any input image, Stable Diffusion first transforms it into a latent representation $x$ using the encoder $\varepsilon$ of a variant auto-encoder  [37]. For any input image, Stable Diffusion first transforms it into a latent representation $x$ using the encoder $\varepsilon$ of a variant auto-encoder  [37]. The diffusion process then operates on $x$ by incrementally introducing noise, resulting in a fixed-length Markov chain represented as $x_{1},x_{2},...,x_{T}$, where $T$ is the chain’s length. Stable Diffusion uses a UNet architecture to learn the reverse of this diffusion process, predicting a denoised version of the latent input $x_{t}$ at each timestep $t$ from 1 to $T$. In the context of text-to-image generation, the text prompts’ conditioning information $y$ is encoded into an intermediate representation $\tau_{\theta}(y)=c$, where $\tau_{\theta}$ is a pre-trained CLIP  [61] text encoder. The primary objective in training this text-to-image diffusion model involves optimizing this transformation and prediction process, and it can be expressed as:
@@
Subject-driven Diffusion Model
Although text-to-image models have achieved remarkable performance, their controllability is limited. To personalize the generated outputs, DreamBooth[66] fine-tunes the diffusion U-Net to fit several target concept images. Custom Diffusion[38] introduces a new modifier token V* in front of the category name and optimizes only the key and value matrices in the cross-attention layers, thereby improving efficiency.
@@ 3.2 Experimental Analysis
We begin by observing simple experimental test cases to realize that the loss of compositional ability after personalization tuning is a common phenomenon. We then analyze the underlying logic through visualizations of the CLIP text-space and cross-attention strength map. Finally, we conduct a theoretical analysis to support our hypothesis.
Simple experimental test cases. As shown in Fig. 1, we observe a loss of compositional ability after fine-tuning. We then conduct additional test cases and find that the headphone concept is not the only one affected; other concepts also experience a similar loss. Furthermore, this situation occurs in both the dog case and other classes.
(a) Comparison of distances in CLIP text space.
(b) Visualization of the cross-attention map activation area.
Figure 3: (a) Each dot represents the position of a phrase combining an adjective and "dog" in the CLIP text-space. After fine-tuning, customized concepts move further away from the the distribution of super-class. (b) Visualization results of cross-attention map activation maps corresponding to the dog token. The bar chart on the right shows the average activation level in the dog area. Experiments show that the activation strengths of the corresponding classes decrease with the increase of the learning rate and the total number of training steps. These demonstrate that the customized concepts likely no longer belong to the super-class, resulting in a loss of super-class semantic information, such as wearing a headphone.
Semantic drift in CLIP text-space. To elucidate the reasons, we project text into the CLIP text space and use two dimensions for simplified visualization in Fig. 3. Each dot represents phrases composing the super-class (“dog” in this case) with different adjectives (e.g., "a cute dog", "a white dog" etc.). Before fine-tuning, the customized concepts (<sks> dog) have no special meaning and are at a similar distance from the super-class as other words. After fine-tuning, we observe a significant increase in the distance of the target concept from its corresponding super-class. This indicates that the semantics of the personalized concepts change during fine-tuning. In short, the model increasingly fails to recognize that personalized concepts belong to the dog category. This shift may lead to an inability to access the knowledge associated with the super-class(like wearing a headphone). More details can be found in Appendix D,  12.
Reduction of cross-attention activation strength. We further investigate the model by visualizing the cross-attention layers in Fig. 3. The attention maps indicate the activation area of super-class words in cross-attention layers. It shows that while the "dog" token activates the relevant region in the image, its activation level is notably lower than that of the pre-trained model. Furthermore, its activation level decreases with the increase in epochs and the learning rate. These findings align with our observations in the CLIP text space and provide support for hypothesis that the customized concepts are increasingly not recognized as part of the super-class during fine-tuning.
Next, we theoretically analyze why the semantic drift of personalized phrases results in the weakening of the compositional ability from the respective reduction in the entropy of composable conditional probability.
@@ 3.3 Theoretical Analysis
$p(x|c_{class},c_{1},c_{2},\dots,c_{i})\propto p(c_{class}|x)p(x)\prod_{i\in T}\frac{p(c_{i})p(x|c_{i})}{p(x)}$ (4)
@@
Figure 4: The orange and green point sets represent the distributions of dogs and headphones, respectively, and their overlapping regions represent their joint probability distributions. During the tuning process, the conditional distribution of dogs and headphones shrinks, which gradually increases the difficulty of sampling. Unlike the Prior Preservation Loss (PPL) in DreamBooth [66], which aims to maintain class diversity, our proposed Semantic Preservation Loss (SPL) focuses on recovering the semantic space of the customized concept. This approach enables our method to synthesize images that are more consistent with the text prompt.
@@
As a result, it is more difficult to sample from our demanded conditional distributions under $c_{\text{class}},c_{1},...,c_{i}$ conditions than before the fine-tuning, leading to the phenomenon that the combining ability is weakened after the fine-tuning. We will discuss the theoretical reasons here in more detail in Appendix B. Fig.  4 illustrates the changes in the distribution during this process that lead to a weakening of the compositional generation capability.
The diminished combinatorial capacity is due to the increased difficulty in sampling from joint conditional probabilities caused by shifts in the semantics of customisation concepts. Therefore, in order to reduce the difficulty of sampling in joint conditional probability distributions, we need to recover the original semantic space of the text, i.e. to recover the semantic distance between custom concepts and superclasses in the semantic space that has been distanced by the fine-tuning process. So, in the next section, we present our proposed semantic preservation loss to mitigate the semantic drift that occurred during fine-tuning.